What is the value of x in the geometric sequence x,2,−12,… ?
Answer by stanbon(75887)

What is the value of x in the geometric sequence x,2,−12,… ?
-----------------
r = -12/2 = -6
-----
x = 2/r
----------------
x = 2/(-6) = -1/3
----------------------
Cheers,
Stan H.
----------------